The manager of a coffee shop has one type of coffee that sells for $5 per pound and one type that sells for $10 per pound. The manager wishes to mix 60 pounds of the $10 coffee to get a mixture that will sell for $8 per pound. How many pounds of the $5 coffee should be used?
Answer by stanbon(75887)

The manager of a coffee shop has one type of coffee that sells for $5 per pound and one type that sells for $10 per pound. The manager wishes to mix 60 pounds of the $10 coffee to get a mixture that will sell for $8 per pound. How many pounds of the $5 coffee should be used?
------------
Equation:
value + value = value
10*60 + 5x = 8(60+x)
600 + 5x = 480 + 8x
3x = 120
x = 40 lbs (amt of $5 coffee needed.
====================================
Cheers,
Stan H.
===============